(3) Using the identity: (*) – 16–191 n! k!(n-k! k for n > 2, prove the following identity: (n-2 + (5+1) 1
Answer #1

We have in i(x-wit We wand to prove .nl () + (+= + first ni (me! (1-2)! (n -(0-2))! íu (1-2)!2! n.(n-1). (1) (ner)! ! =n(n-1)at ut au U 2 an² n २ 60*2)+60*)= 2 o + 11 CS Scanned with CamScanner